(a). (3 points) Suppose the solutions of differential equation xy'''−y'' = 0 are in the form of xr where r is some number. Find three solutions in the form of xr.
(b). (5 points) Find the general solution of xy'''−y'' = 6x^3
Answer #1

xy - 6x3 260 234-De’y - 6% 010-1) (0-2) - 010-) = 0 0² 30²+20 - 070=0 03-40² +30 = 0 o(0-1) (0-3)=0 0=0,1,3 Y = George &